Solve the equation (w-3)^2 minus 36 = 0 where w is a real number. Simplify your answer as much as possible.
Answer by checkley71(8403)

9w-3)~2-36=0 or w~2-6w+9-36=0 or w~2-6w-27 or (w-9)(w+3)=0 or w-9=0 or w=9 &
w+3=0 or w=-3